Decide whether the number is a solution of the equation.
| 2x+3 | =7;2 I am thinking that if x =2, the solution is 7. Am I on the right track? Thank you.

Answer by Alan3354(69443)

| 2x+3 | =7;2
----------
2x+3 = 7 --> x = 2
-----
-(2x+3) = 7
x = -5
----------
2 solutions, 2 & -5